OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 13, c = 21 15.83 15.87 16.19 16.49

OpenStudy (anonymous):

quadratics to triangles in one step must be an on line class....

OpenStudy (anonymous):

16.49?

OpenStudy (anonymous):

\[a^2+b^2=c^2\] is the conclusion of pythagoras the one theorem everyone knows so \[a^2+13^2=21^2\] you gotta solve for \(a\)

OpenStudy (anonymous):

You have to use the Pythagorean theorem

OpenStudy (anonymous):

yeah you got it

OpenStudy (anonymous):

thanks